Scheduling Meeting Minutes: 14 Jan 2008
Monday afternoon: Access request by RF to study Stochastic Cooling Status.
Ten minute access requested to check whether repairs are possible in 5-10
minutes. If not, then pull the power from the motors that drive the
kickers. This also gives the opportunity to prepare for any serious
repairs on Wednesday. If work is needed on the SC on Wednesday, then the
work scheduled on the inchworm on Wednesday would be bumped for SC repiar.
While the Stochastic Cooling work is going on, STAR will take the opportunity
of an access to take care of 3 small jobs inside. No access request for
PHENIX.
Tuesday: APEX starts 8PM
Wednesday: Maintenance starts 8AM
Polarized Proton Turn On
This schedule assumes:
•
No 9 MHz RF in RHIC.
•
No WF3 repair.
•
No AGS Jump Target pin diodes.
•
Rotators for PHENIX only.
Starting Monday January 7th
•
LINAC startup for MODS 6-9. (Done?)
•
Polarized source work.
•
Pre-testing of AGS pp quads and bumps.
•
Pre-testing of RHIC snake and rotator power supplies.
•
Continue cool down of AGS cold snake.
•
RSC Checklists for protons in LTB, Booster AGS and RHIC.
•
Live testing of snake correction quads and bumps behind RHIC stores.
Monday January 14th
• AM:
Helium introduced into A20 Snake in AGS.
•
PM: Hi Pot A20 snake from A18 house.
Tuesday January 15th:
•
Test AGS cold snake with current (behind RHIC store)
Maintenance additions to January 16
•
AGS C20 polarimeter work.
•
HEBT Vacuum work (NO BLIP for 10-12 hours).
•
Solenoid installation in LEBT.
•
200 MeV polarimeter work.
•
RHIC rotator and snake start up.
•
RHIC polarimeter testing/work. (APEX?)
•
Final check out bumps and quads for pp in AGS.
•
PP2PP access for instrumentation, experimental, video, mechanical and Access
Controls.(?)
•
JET Outlet get installed.
Upon completion of the d-Au run
Access to RHIC for:
•
Power supply work-
–Q7
polarity reversal,
–snake
rotator final checkout.
•
Move ZDCs at STAR and PHENIX to PP position (ES&F)
Install scintillators and survey.
•
Booster flags at extraction (1/2 day).
•
Tentative: AGS C15 Polarimeter work
For Au31+
in AGS:
•
The A5 kicker magnet will be running near max output.
•
Modification of the ion pump readout system will be complete and tested by
January 18th,
2008 at the latest.
•
4 hour Sextupole polarity swap in RHIC – if this is done last, would not
have to be reversed until shutdown.
Polarized proton equipment readiness list
•
Quads: A19, B1, B3 and E17.
•
Snakes I10 – contact S. Savatteri prior to energizing.
•
E20 Snake
•
Correction bump A16-A19
•
Correction bump B2-B5
•
LINAC Mods for Cavities 6-9
•
Hi-pot and testing of AGS cold snake
Polarimeter
JET work for Monday January 28th
is
to replace pump 5-2:
•
bleeding up the beam region vacuum (1/2 hr),
•replace
the pump (2 hrs),
•pump
down to some reasonable state (1 hr
•re-activate
the NEG cartridge (4hrs).
•turn
on the jet dissociator (3-4 hrs for cooling down and tuning up).
The shutdown for JET will require 12 hours total.
Carbon Polarimeter
•
RHIC Blue should be battle-tested tomorrow
•
RHIC Yellow should be ready after Maintenance Day 16 Jan preamp installation
•
AGS C15 should be ready
•
AGS C20 (?)
PHENIX and STAR Experiment Work
•
No specific requests from either experiment for work that is required to
transition from d-Au to pp
•PHENIX
wants to purge their Pad Chamber gas system with alcohol, and that may take
several hours of access.
PP2PP
Install silicon packages on 28 Jan (12 hours)
More detailed schedule will be supplied after this week's maintenance day.
Scheduling Physicist: Mike Sivertz
| Last updated:
|
Email:
sivertz@bnl.gov |
Office Phone: 631-344-6102 |
Digital Pager: 7528 or 4300 |
Home Phone: 631-878-9020 (leave message) |